A 21.0 g ball is fired horizontally with initial speed v0 toward a 110 g ball that is hanging motionless from a 1.10 m -long string. The balls undergo a head-on, perfectly elastic collision, after which the 110 g ball swings out to a maximum angle θmax = 52.0

What was v0? Express your answer with the appropriate units.
Answer #1

Height, h = L(1 - cos(x)) = 1.1(1 - cos 52)) =0.423 m

Velocity, v2 = sqrt (2 x 9.8 x 0.423) = 2.88 m/s

2 Vo = V2 (1 + m2/m1)

Vo = 0.5 x 2.88 (1 + 110/21) = 8.98 m/s

Comment in case any doubt please rate my answer... .